OpenStudy (anonymous):

A ball of radius 13 has a round hole of radius 6 drilled through its center. Find the volume of the resulting solid.

OpenStudy (anonymous):

What do I use to find the volum of the second structure?

OpenStudy (anonymous):

oh so pi*r^2*h

OpenStudy (anonymous):

Pi*12^2*26

OpenStudy (anonymous):

And 4/3 pi r^3 for the ball

OpenStudy (anonymous):

(4/3) pi 6^3

OpenStudy (anonymous):

that is of this cone

OpenStudy (zarkon):

now I don't know what you are doing :)

OpenStudy (anonymous):

1/3 pi r^2 h so (1/3)pi 12^2 * 13

OpenStudy (zarkon):

I would start with \[x^2+y^2+z^2=13^2\] then \[z=\pm\sqrt{13^2-x^2-y^2}\] then I would integrate between these two function over the region defined by the hole. then subtract this volume from the volume of the sphere

OpenStudy (zarkon):

to integrate one should switch to polar coordinates... I get \[\frac{532\pi\sqrt{133}}{3}\]

OpenStudy (anonymous):

That worked! Thanks a lot!

OpenStudy (zarkon):

just so everyone can see this is what I did... \[\frac{4}{3}\pi 13^3-2\int\limits_{0}^{2\pi}\int\limits_{0}^{6}\sqrt{13^2-r^2}r\,dr\,d\theta\]